The marks of 500 candidates in an examination are normally distributed with a mean of 45 marks and a standard deviation of 20 marks. Given that the pass marks is 41, estimate the of candidates, who passed the examination


1
Expert's answer
2021-11-03T13:49:44-0400

"\mu=45 \\

\sigma=20 \\

P(X>41) = 1 -P(X<41) \\

= 1 -P(Z < \frac{41-45}{20}) \\

= 1 -P(Z< -0.2) \\

= 1 -0.4207 \\

= 0.5793"

Candidates, who passed the examination "= 500 \times 0.5793 = 289.65 ≈ 290"